# When, Why and How Much? Adaptive Learning Rate Scheduling by Refinement

### v1

@@ 5.3 Any-time Adaptive Schedules
Methods based on hyper-gradients (i.e. gradients of the learning rate) are another heavily studied class of any-time adaptive methods. They use recent gradient evaluations and other local information to update the learning rate at each step of training (Almeida et al., 1999; Franceschi et al., 2017; Bengio, 2000; Domke, 2012; Pedregosa, 2016; Baydin et al., 2018; Feurer and Hutter, 2019; Donini et al., 2020; Chandra et al., 2022). The hyper-gradient can also be used in a more cautious manor to define conditions under which the learning rate should be decreased (Pflug, 1983, 1988; Lang et al., 2019; Zhang et al., 2020).
Methods based on loss minimization aim to choose a learning rate that greedily minimize the loss, either at each step, or based on aggregate statistics that are either produced in stages or by smoothing. They include methods that target the train loss or loss on held out validation data (Xu et al., 2019; Jin et al., 2021; Kim et al., 2021).
Methods that query additional points, such as line search methods, are not scheduling methods per se but can be used in place of scheduling (Mahsereci and Hennig, 2017; Iyer et al., 2021). Due to the additional implementation complexity and the cost of their use, they also have not seen widespread adoption for deep learning.
Conclusion
Our experimental results strongly support the use of the linear decay schedule as a default baseline for deep learning optimization. Applying refinement on top of an initial linear decay run gives a schedule showing improvements on 5 out of 10 deep learning problems and 5 of 8 convex problems. We develop extensive theory supporting the linear decay schedule and our novel refinement technique, establishing that their SOTA performance is well-supported both in theory and practice.
Appendix A All-Tail Summation Bound
@@
In this section we develop some key Lemmas that underlie our results. The starting point of this development is the following result from Orabona (2020); Lin et al. (2016). We present the result and proof below for completeness, and then provide an improved version of the bound (Lemma 5) that is used to derive our analytical results.
A.1 Existing bound
Lemma 4
@@ A.1 Existing bound
Proof  We present below a version of the proof that’s written to move the single inequality application in the whole proof to the last step. Define $S_{k}=\frac{1}{k}\sum_{t=T-k+1}^{T}\eta_{t}q_{t}$.
@@
Where the final step uses the fact that $\eta_{t}$ is decreasing. Plugging in the definition of $S_{T}$, and substituting $k=T-k$ to simplify gives the result.
A.2 Improved expression
Lemma 5
@@ A.2 Improved expression
Finally, we make the simple change of variables $k\to T-k$ to yield the result.
Remark 6
@@
Appendix B Proof of Theorem  1
@@
Suppose $g_{t}$ are random variables. with $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ for some convex $f$. Then:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-u\rangle\right].$
Proof  Let us set $q_{t}=f(x_{t})-f(u)$. Then we have $\mathbb{E}[q_{t}]\leq\mathbb{E}[\langle g_{t},x_{t}-u\rangle]$ and $\mathbb{E}[q_{t}-q_{k}]\leq\mathbb{E}[\langle g_{t},x_{t}-x_{k}\rangle]$. Then, Lemma 5 implies:
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},x_{t}-u\rangle+\sum_{k=1}^{T-1}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)\sum_{t=k}^{T}w_{t}\langle g_{t},x_{t}-x_{k}\rangle\right].$
@@
And for $p>t$, the coefficient of $\langle g_{t},x_{p}\rangle$ is zero. Finally the coefficient of $\langle g_{t},u\rangle$ is $-\frac{w_{t}}{w_{1:T}}$.
@@
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle g_{t},\frac{w_{t}}{w_{t:T}}x_{t}-\left(\frac{w_{t}u}{w_{1:T}}+\sum_{p=1}^{t-1}w_{t}x_{t}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{u}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right].$
@@
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{u}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\frac{w_{t}}{w_{1:T}}\langle g_{t},z_{t}-u\rangle\right].$
Now, we are finally ready to prove Theorem 1:
See 1
Proof  Let’s define $\hat{x}_{1}=z_{t}$ and recursively set:
@@
Then, Lemma 7 shows that $\mathbb{E}[f(\hat{x}_{T})-f(u)]\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-u\rangle\right]$. So, it suffices to show that $x_{t}=\hat{x}_{t}$ for all $t$. In turn, since $\hat{x}_{1}=z_{1}=x_{1}$, it suffices to show $\hat{x}_{t+1}-\hat{x}_{t}=\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}=x_{t+1}-x_{t}$ for all $t$.
@@
Appendix C Proof of Theorem 3
@@
Proof  First, observe that with $\Delta_{t}=-w_{t}\eta_{t}$, we have $x_{t+1}=x_{t}+\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}$. Therefore with $z_{1}=x_{1}$ and $z_{t+1}=z_{t}+\Delta_{t}$, Theorem 1 implies:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle\right].$
Next, observe that $z_{t}$ is simply online gradient descent with learning rate 1 acting on the loss vectors $w_{t}z_{t}$. Standard analysis (Zinkevich, 2003) shows:
$\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle$ $=\frac{\|z_{1}-u\|^{2}}{2}-\frac{\|z_{T+1}-u\|^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}\|g_{t}\|^{2}}{2}.$
This immediately implies the first part of the Theorem. Next, we need to solve for the minimizing values of $w_{t}$. To do this, we take the logarithm of the expression $\frac{1}{2w_{1:T}}\left(\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)$ and differentiate:
$\frac{\partial}{\partial w_{k}}\log\left[\frac{1}{2w_{1:T}}\left(\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right]$ $=\frac{2w_{k}\|g_{k}\|^{2}}{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}-\frac{1}{w_{1:T}}.$
@@
$w_{k}$ $=\|g_{k}\|^{-2}\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}\triangleq\lambda\|g_{k}\|^{-2},$
where we have defined $\lambda=\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$, which does not depend on $k$. That is, the optimal $w_{k}$ value is proportional to $\|g_{k}\|^{-2}$. With this expression, we have:
@@
$\lambda$ $=\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$
$\lambda$ $=\frac{\|x_{1}-u\|^{2}+\lambda^{2}\sum_{t=1}^{T}\|g_{t}\|^{-2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}$
$\lambda$ $=\frac{\|x_{1}-u\|^{2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}+\frac{\lambda}{2}$
$\lambda$ $=\frac{\|x_{1}-u\|}{\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{-2}}}.$
This in turn implies the claimed optimal value for $w_{k}$.
Appendix D Proof of Theorem 9
Proof  First, observe that for any solution to the desired identity $\eta_{t}=\frac{w_{t}w_{t:1:T}}{w_{1:T}}$, replacing $w$ with $c\cdot w$ for some constant $c$ will yield a solution for $\eta$ replaced with $c\cdot\eta$. Thus, it suffices to consider the possibility that $\max_{t}\eta_{t}=1$.
@@ Appendix D Proof of Theorem 9
Next, to show (2): $s_{t}=w_{t:T}$. This is nearly immediate. By definition we have $w_{T}=s_{T}$. Suppose $s_{k}=w_{t:T}$ for some $t\geq 2$. Then $s_{t}=s_{t-1}-w_{t-1}$ so that $s_{t-1}=s_{t}+w_{t-1}=w_{t-1:T}$. Thus by induction we have $s_{t}=w_{t:T}$ for all $t$, which is the last thing we needed to show.
Appendix E Schedules for Per-Coordinate Updates
Many popular optimization algorithms in use today like Adam (Kingma and Ba, 2015) and its variants employ per-coordinate updates: the update $\Delta_{t}$ is not proportional to the gradient $g_{t}$ but instead scales each coordinate of $g_{t}$ by an adaptively chosen value. In this section we propose an approximately optimal schedule for such methods.
Theorem 8
$R=\sqrt{\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}-(z_{t+1,i}-u_{i})^{2}}{\eta_{t,i}}}$
Then we have:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\left(\frac{R^{2}}{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)\right].$
$w_{t}=\lambda\left(\sum_{i=1}^{d}\frac{g_{t,i}^{2}}{\sqrt{v_{t,i}}}\right)^{-1}$
[…]
3:Define $w_{t}=\hat{G}_{t}^{-1}$
4:For each $t$, let:
5:
$\eta_{t}=w_{t}\sum_{p=t+1}^{T}w_{p}$
6:Return normalized schedule $\eta/\max(\eta)$
Algorithm 2 Schedule Refinement for Adam
Proof [proof of Theorem 8] By standard online gradient descent analysis, we have:
$\langle w_{t}\cdot g_{t},z_{t}-u\rangle$ $=\sum_{i=1}^{d}w_{t}g_{t,i}(z_{t,i}-u_{i})$
$=\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-u_{i})^{2}}{2\eta_{t,i}}+\frac{\eta_{t,i}}{2}w_{t}^{2}g_{t,i}^{2}.$
Summing this over $t$ from $1$ to $T$, we get
$\sum_{t=1}^{T}\langle w_{t}\cdot g_{t},z_{t}-u\rangle$ $=\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-u_{i})^{2}}{2\eta_{t,i}}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\frac{\eta_{t,i}g_{t,i}^{2}}{2}$
$=\frac{R^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}}{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}.$
The first part of the result now follows from Theorem 1.
Rearranging, we obtain
$\triangleq\lambda\left(\sum_{i=1}^{d}\eta_{k,i}g_{k,i}^{2}\right)^{-1},$
Putting $w_{t}=\lambda\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}$ completes the proof.
Appendix F Explicit bounds for arbitrary schedules
Theorem 9
@@ Appendix F Explicit bounds for arbitrary schedules
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{1}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right]$
@@
This can be bounded by the usual suboptimality inequality for SGD/GD methods, which holds for any $u$:
$\sum_{t=k}^{T}\eta_{t}\left[f(x_{t})-f(u)\right]\leq\frac{1}{2}\left\|x_{k}-u\right\|^{2}+\frac{1}{2}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2},$
Setting $u=x_{k}$ yields:
@@
Now, we are ready to prove Theorem 10. See 10
Proof  Following the standard convergence bound approach:
$\left\|x_{t+1}-u\right\|^{2}$ $=\left\|x_{t}-\eta_{t}g_{t}-u\right\|^{2}$
$=\left\|x_{t}-u\right\|^{2}-2\eta_{t}\left\langle g_{t},x_{t}-u\right\rangle+\eta_{t}^{2}\left\|g_{t}\right\|^{2}$
$\leq\left\|x_{t}-u\right\|^{2}-2\eta_{t}\left[f(x_{t})-f(u)\right]+\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$

### v2

@@ 5 Related Work
Methods based on hyper-gradients (i.e. gradients of the learning rate) are another well-studied class of adaptive method. They use recent gradient evaluations and other local information to update the learning rate (Almeida et al.,, 1999; Franceschi et al.,, 2017; Bengio,, 2000; Domke,, 2012; Pedregosa,, 2016; Baydin et al.,, 2018; Feurer and Hutter,, 2019; Donini et al.,, 2020; Chandra et al.,, 2022). Alternatively, hyper-gradients can be used to signal when to decrease the learning rate (Pflug,, 1983, 1988; Lang et al.,, 2019; Zhang et al.,, 2020). Methods based on loss minimization choose a learning rate that greedily minimizes the test or train loss (Xu et al.,, 2019; Jin et al.,, 2021; Kim et al.,, 2021).
Appendix A Limitations
There are certain problems like CIFAR-10 and CIFAR-100 where we can clearly observe potential pitfalls of our approach (Figure 5) when our assumptions break down. On these problems $100\%$ train accuracy is reached near the end, driving the gradient norm sequence to zero. The resulting refined schedule rapidly increases the step size near the end of optimization, which if used results in training divergence. We suggest using a linear decay schedule instead on problems where the gradient norm drops significantly at the end.
Our work assumes that multiple training runs for the given workload are possible. This often holds in practice, particularly in industry settings where the same models are repeatedly trained on fresh data. Even in cases where large one-off training runs are performed, refined schedules can be extrapolated from smaller trial runs, or similar past runs.
Although our theory is limited to the convex setting, it seems to have predictive power for non-convex problems also. The goal of our theory is to guide practice, not to provide an explanation of potential pathological worst case behavior that can occur in non-convex settings. For this purpose, simplified settings are ideal as they allow for precise and predictive results.
Appendix B All-Tail Summation Bound
@@
In this section we develop some key Lemmas that underlie our results. The starting point of this development is the following result from Orabona, (2020); Lin et al., (2016). We present the result and proof below for completeness, and then provide an improved version of the bound (Lemma 5) that is used to derive our analytical results.
B.1 Existing bound
Lemma 4.
@@ B.1 Existing bound
Proof.
We present below a version of the proof that’s written to move the single inequality application in the whole proof to the last step. Define $S_{k}=\frac{1}{k}\sum_{t=T-k+1}^{T}\eta_{t}q_{t}$.
@@
Where the final step uses the fact that $\eta_{t}$ is decreasing. Plugging in the definition of $S_{T}$, and substituting $k=T-k$ to simplify gives the result. ∎
B.2 Improved expression
Lemma 5.
@@ B.2 Improved expression
Finally, we make the simple change of variables $k\to T-k$ to yield the result. ∎
Remark 6.
@@
Appendix C Proof of Theorem  1
@@
Suppose $g_{t}$ are random variables with $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ for some convex $f$. Then:
$\mathbb{E}[f(x_{T})-f(x_{*})]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-x_{*}\rangle\right].$
Proof.
Let us set $q_{t}=f(x_{t})-f(x_{*})$. Then we have $\mathbb{E}[q_{t}]\leq\mathbb{E}[\langle g_{t},x_{t}-x_{*}\rangle]$ and $\mathbb{E}[q_{t}-q_{k}]\leq\mathbb{E}[\langle g_{t},x_{t}-x_{k}\rangle]$. Then, Lemma 5 implies:
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},x_{t}-x_{*}\rangle+\sum_{k=1}^{T-1}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)\sum_{t=k}^{T}w_{t}\langle g_{t},x_{t}-x_{k}\rangle\right].$
@@
And for $p>t$, the coefficient of $\langle g_{t},x_{p}\rangle$ is zero. Finally the coefficient of $\langle g_{t},x_{*}\rangle$ is $-\frac{w_{t}}{w_{1:T}}$.
@@
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle g_{t},\frac{w_{t}}{w_{t:T}}x_{t}-\left(\frac{w_{t}x_{*}}{w_{1:T}}+\sum_{p=1}^{t-1}w_{t}x_{t}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{x_{*}}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right].$
@@
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{x_{*}}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\frac{w_{t}}{w_{1:T}}\langle g_{t},z_{t}-x_{*}\rangle\right].$
Now, we are finally ready to prove Theorem 1: See 1
Proof.
Let’s define $\hat{x}_{1}=z_{t}$ and recursively set:
@@
Then, Lemma 7 shows that $\mathbb{E}[f(\hat{x}_{T})-f(x_{*})]\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-x_{*}\rangle\right]$. So, it suffices to show that $x_{t}=\hat{x}_{t}$ for all $t$. In turn, since $\hat{x}_{1}=z_{1}=x_{1}$, it suffices to show $\hat{x}_{t+1}-\hat{x}_{t}=\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}=x_{t+1}-x_{t}$ for all $t$.
@@
Appendix D Proof of Theorem 3
@@
Proof.
First, observe that with $\Delta_{t}=-w_{t}g_{t}$, we have $x_{t+1}=x_{t}+\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}$. Therefore with $z_{1}=x_{1}$ and $z_{t+1}=z_{t}+\Delta_{t}$, Theorem 1 implies:
$\mathbb{E}[f(x_{T})-f(x_{*})]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-x_{*}\rangle\right].$
Next, observe that $z_{t}$ is simply online gradient descent with learning rate 1 acting on the loss vectors $w_{t}z_{t}$. Standard analysis (Zinkevich,, 2003) shows:
$\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-x_{*}\rangle$ $=\frac{\|z_{1}-x_{*}\|^{2}}{2}-\frac{\|z_{T+1}-x_{*}\|^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}\|g_{t}\|^{2}}{2}.$
This immediately implies the first part of the Theorem. Next, we need to solve for the minimizing values of $w_{t}$. To do this, we take the logarithm of the expression $\frac{1}{2w_{1:T}}\left(\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)$ and differentiate:
$\frac{\partial}{\partial w_{k}}\log\left[\frac{1}{2w_{1:T}}\left(\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right]$ $=\frac{2w_{k}\|g_{k}\|^{2}}{\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}-\frac{1}{w_{1:T}}.$
@@
$w_{k}$ $=\|g_{k}\|^{-2}\frac{\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}\triangleq\lambda\|g_{k}\|^{-2},$
where we have defined $\lambda=\frac{\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$, which does not depend on $k$. That is, the optimal $w_{k}$ value is proportional to $\|g_{k}\|^{-2}$. With this expression, we have:
@@
$\lambda$ $=\frac{\|x_{1}-x_{*}\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$
$\lambda$ $=\frac{\|x_{1}-x_{*}\|^{2}+\lambda^{2}\sum_{t=1}^{T}\|g_{t}\|^{-2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}$
$\lambda$ $=\frac{\|x_{1}-x_{*}\|^{2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}+\frac{\lambda}{2}$
$\lambda$ $=\frac{\|x_{1}-x_{*}\|}{\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{-2}}}.$
This in turn implies the claimed optimal value for $w_{k}$. ∎
Appendix E Schedules for Per-Coordinate Updates
Many popular optimization algorithms in use today like Adam (Kingma and Ba,, 2015) and its variants employ per-coordinate updates: the update $\Delta_{t}$ is not proportional to the gradient $g_{t}$ but instead scales each coordinate of $g_{t}$ by an adaptively chosen value. In this section we propose an approximately optimal schedule for such methods.
Theorem 8.
$R=\sqrt{\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-{x_{*}}_{i})^{2}-(z_{t+1,i}-{x_{*}}_{i})^{2}}{\eta_{t,i}}}$
Then we have:
$\mathbb{E}[f(x_{T})-f(x_{*})]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\left(\frac{R^{2}}{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)\right].$
$w_{t}=\lambda\left(\sum_{i=1}^{d}\frac{g_{t,i}^{2}}{\sqrt{v_{t,i}}}\right)^{-1}$
2:  $\hat{G}=\textrm{median\_filter}(G,\textrm{filter\_width}=\tau\,T,\textrm{padding}=(\textrm{nearest},\textrm{reflect}))$
3:  Define $w_{t}=\hat{G}_{t}^{-1}$
4:  For each $t$, let:
5:
$\eta_{t}=w_{t}\sum_{p=t+1}^{T}w_{p}$
6:  Return normalized schedule $\eta/\max(\eta)$
Algorithm 2 Schedule Refinement for Adam
proof of Theorem 8.
By standard online gradient descent analysis, we have:
$\langle w_{t}\cdot g_{t},z_{t}-x_{*}\rangle$ $=\sum_{i=1}^{d}w_{t}g_{t,i}(z_{t,i}-{x_{*}}_{i})$
$=\sum_{i=1}^{d}\frac{(z_{t,i}-{x_{*}}_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-{x_{*}}_{i})^{2}}{2\eta_{t,i}}+\frac{\eta_{t,i}}{2}w_{t}^{2}g_{t,i}^{2}.$
Summing this over $t$ from $1$ to $T$, we get
$\sum_{t=1}^{T}\langle w_{t}\cdot g_{t},z_{t}-x_{*}\rangle$ $=\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-{x_{*}}_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-{x_{*}}_{i})^{2}}{2\eta_{t,i}}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\frac{\eta_{t,i}g_{t,i}^{2}}{2}$
$=\frac{R^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}}{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}.$
The first part of the result now follows from Theorem 1.
Rearranging, we obtain
$\triangleq\lambda\left(\sum_{i=1}^{d}\eta_{k,i}g_{k,i}^{2}\right)^{-1},$
Putting $w_{t}=\lambda\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}$ completes the proof. ∎
Appendix F Explicit Bounds for Arbitrary Schedules
Theorem 9.
Proof.
First, observe that for any solution to the desired identity $\eta_{t}=\frac{w_{t}w_{t:1:T}}{w_{1:T}}$, replacing $w$ with $c\cdot w$ for some constant $c$ will yield a solution for $\eta$ replaced with $c\cdot\eta$. Thus, it suffices to consider the possibility that $\max_{t}\eta_{t}=1$.
@@ Appendix F Explicit Bounds for Arbitrary Schedules
Next, to show (2): $s_{t}=w_{t:T}$. This is nearly immediate. By definition we have $w_{T}=s_{T}$. Suppose $s_{k}=w_{t:T}$ for some $t\geq 2$. Then $s_{t}=s_{t-1}-w_{t-1}$ so that $s_{t-1}=s_{t}+w_{t-1}=w_{t-1:T}$. Thus by induction we have $s_{t}=w_{t:T}$ for all $t$, which is the last thing we needed to show. ∎
@@
$\mathbb{E}[f(x_{T})-f(x_{*})]$ $\leq\mathbb{E}\left[\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{1}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right]$
@@
This can be bounded by the usual suboptimality inequality for SGD/GD methods, which holds for any $x_{*}$:
$\sum_{t=k}^{T}\eta_{t}\left[f(x_{t})-f(x_{*})\right]\leq\frac{1}{2}\left\|x_{k}-x_{*}\right\|^{2}+\frac{1}{2}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2},$
Setting $x_{*}=x_{k}$ yields:
@@
F.1 Proof of Theorem 10
Proof.
Following the standard convergence bound approach:
$\left\|x_{t+1}-x_{*}\right\|^{2}$ $=\left\|x_{t}-\eta_{t}g_{t}-x_{*}\right\|^{2}$
$=\left\|x_{t}-x_{*}\right\|^{2}-2\eta_{t}\left\langle g_{t},x_{t}-x_{*}\right\rangle+\eta_{t}^{2}\left\|g_{t}\right\|^{2}$
$\leq\left\|x_{t}-x_{*}\right\|^{2}-2\eta_{t}\left[f(x_{t})-f(x_{*})\right]+\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$